# Shift current response in twisted double bilayer graphenes

### v1

@@ III.1 Intrinsic shift current response
The results for the $\sigma^{x}_{xx}$ response in AB-AB and AB-BA stacked TDBG are shown on the leftmost panels in Fig. 2 (a) and (b) respectively. Focussing on the AB-AB variant, we show the band structure on the left panel and $\sigma^{x}_{xx}(\omega;E_{\rm{F}})$ as a colour density plot on the right panel. The shift current is plotted with frequency $\omega=0\sim 70$ meV on the horizontal and Fermi level $E_{\rm{F}}=-50\sim 50$ meV on the vertical axis. Therefore, each horizontal strip of the density plot corresponds to a $\sigma^{x}_{xx}$ $\omega$ plot for a certain value of $E_{\rm{F}}$. We note that the scale of the vertical axis is adjusted to match that of the band structure plot. In Fig. 2 (b), we show the corresponding plots for AB-BA TDBG, however, due the the $C_{2y}$ symmetry, $\sigma^{x}_{xx}$ is completely suppressed.
If we inspect the $\sigma^{x}_{xx}$ density plot in AB-AB TDBG, we see that there is a strong signal in the low frequency region which gradually tails off as we move to higher frequencies. This behaviour is seen throughout the plotted $E_{\rm{F}}$ range and is due to the $1/\omega^{2}$ prefactor in Eq. (6). This prefactor acts to enhance the signal when the gap is small and to cause a suppression for high frequencies. The enhancement is especially strong near the charge neutrality point (CNP), indicated by the grey arrow, where bands are concentrated giving rise to very small energy gaps. A similar enhancement of the shift current is reported other moiré materials [31; 72; 73; 74]. We also notice that the low frequency signal is not of a single sign but a series of sign changes is observed as the Fermi level is swept. Note that the signal should technically vanish as $\omega\to 0$, however, due to to the finite broadening and the colour scale of the density plot, it appears that there is a finite response in the dc limit.
The results for the $\sigma^{y}_{yy}$ response are shown in Fig. 3 in a similar fashion. In this case, the (b) AB-BA variant gives a finite response while it vanishes in the (a) AB-AB case due to the $C_{2x}$ symmetry. Although there are differences between the $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$ responses, we find that their overall behaviour, such as the strong enhancement at low frequencies, are common.
@@ III.2 Effect of vertical bias voltage
Figure 2: The band structure and $\sigma^{x}_{xx}$ density plots of (a) AB-AB and (b) AB-BA stacked TDBG for various strength of the vertical bias voltage $\Delta$. Here, we used $\Delta=$ 0, 10, 20 and 50 meV. The grey arrows indicate the respective charge neutral gaps. We also show the moiré Brillouin zone in the top right corner.
Figure 3: The band structure and $\sigma^{y}_{yy}$ density plots of (a) AB-AB and (b) AB-BA stacked TDBG for various strength of the vertical bias voltage $\Delta$. Here, we used $\Delta=$ 0, 10, 20 and 50 meV. The grey arrows indicate the respective charge neutral gaps.
Having studied the behaviour of the intrinsic shift current response, we now move our attention to the evolution of the response as a result of the application of $\Delta$. We begin with the $\sigma^{x}_{xx}$ response at values of $\Delta=$ 10, 20 and 50 meV and show the band structures of AB-AB and AB-BA stacked TDBG in Fig. 2 (a) and (b) respectively.
Focussing on the AB-AB configuration, where $\sigma^{x}_{xx}$ is always non-zero, as $\Delta$ is increased, we see that the overall size of the signal is suppressed. This is due to the fact that the bias voltage increases the gap size, which in turn suppresses the signal via the $1/\omega^{2}$ factor. On the other hand, in the AB-BA variant, as soon as the vertical bias voltage begins to break the $C_{2y}$ symmetry, we observe a finite response. From here, the signal continues to grow in magnitude until $\Delta\sim$ 10 meV, from which the signal follows the same pattern with the $1/\omega^{2}$ suppression for stronger $\Delta$. A similar pattern between the signal strength and the band gap can be seen when the twist angle $\theta$ is increased. The results for $\theta=0.4^{\circ}$ and $2.0^{\circ}$ at $\Delta=$ 50 meV are given in Appendix A.
@@ III.3 Comparison between AB-AB and AB-BA stacked TDBG
If we begin to compare the density plots between the two stacking configurations, we notice that there is a general positive/negative correlation in the sign of the signal above/below the CNP. The negative correlation is illustrated in Fig. 4 where we show a close-up view of the four density plots below the CNP, indicated by the grey dashed boxes in Fig. 2 and 3. Focussing first on the $\sigma^{x}_{xx}$ response shown in Fig. 4 (a) and (b), we have labelled four regions from A to D where a sign reversal can be seen. For example, in the AB-AB variant, region A has a positive signal, while the corresponding signal in the AB-BA variant is negative. A similar pattern is seen for the other three regions. Likewise, in Fig. 4 (c) and (d), where show the density plots for $\sigma^{y}_{yy}$, we find that regions with the same label have opposite signs. We note that this positive/negative correlation is overshadowed as we turn down $\Delta$ and the bands begin to cluster.
Figure 4: A close-up view of the $\sigma^{x}_{xx}$/$\sigma^{y}_{yy}$ responses in (a)/(c) AB-AB and (b)/(d) AB-BA TDBG shown in Fig. 2 and 3. We have labelled regions where a sign reversal between the two stacking configurations can be seen.
@@ IV Understanding sign reversal in TDBG shift current
Using this Hamiltonian, we show the band structure (left panel) and the corresponding $\sigma^{y}_{yy}$ response (right panel) of BLG in Fig. 5. The band structure computed from the Hamiltonian in Eq. (11) shows a slice through $k_{x}$ at $k_{y}=0$ around the $K_{+}$ ($K_{-}$) point, plotted in the black solid (red dashed) line. The vertical bias voltage allows for a gap-opening of around 50 meV. On the right panel, we show the shift current response as a colour density plot with frequency $\omega$ on the horizontal and Fermi level $E_{\rm{F}}$ on the vertical axis. The numerical evaluation of Eq. (7) was performed with a broadening of $\eta=$ 1 meV. We can initially see that there is very little or no response at frequencies below 50 meV, which is a reflection of the fact that there are no possible optical transitions in this frequency range due to the gap opened via the bias voltage. As we move up to $\omega=$ 50 meV, we start to see a large response when the Fermi level is set within the gap. This sharp response corresponds to the optical transition from the band edge of the valence to that of the conduction band, where a large joint density of states (JDoS) is realised. We also note that the size of the signal is in the range of 10 mA nm $\rm{V}^{-2}$, which is an order of magnitude smaller compared to the signal seen in TDBG. This result suggests that the effect of the moiré reconstruction of the band structure indeed enhances the shift current response.
Figure 5: The band structure of AB-stacked BLG from the $K_{+}$ ($K_{-}$) is plotted in the solid black (dashed red) lines in the left panel; a vertical bias voltage of 50 meV is applied across the layers. The shift current $\sigma^{y}_{yy}$ is plotted as a density plot in the right panel.
@@
where we have used the fact that $\sigma^{x^{\prime}}_{x^{\prime}x^{\prime}}=0$. The detail of the derivation is described in Appendix. B. We note that, now, both $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$ are finite since $M_{x}$ symmetry is broken as the two BLG sheets are twisted relative to each other.
@@
If we focus on the signs of the $\sigma^{x}_{xx}$ signal on the centre panels, we can see that there is a positive/negative correlation above/below the charge neutrality point between the two variants. This can be understood from the fact that the relative orientation of the BLG sheets are aligned/antialigned in the AB-AB/AB-BA variant. As the shift current is related to the polarisation of the material, the $180^{\circ}$ rotation of the second layer leads to a sign reversal of the signal. The same effect is seen in the $\sigma^{y}_{yy}$ signal on the right panels. This qualitatively explains the sign reversal that was seen in the TDBG calculation, and it reveals that this relationship is retained even after the effect of the moiré coupling.
@@
Figure 7: The band structure of (a) AB-AB and (b) AB-BA stacked uncoupled TDBG is shown in the left panels. The plot is around the $K_{+}$ point, the twist angle is set to $\theta=0.8^{\circ}$ and a vertical bias voltage of $\Delta=$ 50 meV is applied. The contribution from the first (second) BLG is plotted in the solid (dotted) black lines. The corresponding moiré reciprocal lattice vector is also labelled. On the centre and right panels, we show the corresponding $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$ plots, plotting frequency $\omega$ and Fermi energy $E_{\rm{F}}$ on the horizontal and vertical axes respectively.
@@ V Conclusion
In the present paper, we have studied the shift current response in AB-AB and AB-BA stacked twisted double bilayer graphene (TDBG). We have studied the intrinsic signal and the effect on the signal as we turn on the vertical bias voltage and their dependence on the Fermi level. In the low frequency regime, we have found a strong signal and a series of sign flips across band edges. We have further found a positive/negative correlation in the sign of the signal between the AB-AB and AB-BA variants above/below the charge neutrality point at large values of the vertical bias voltage. To understand the origin of the sign relationship between the two variants, we used a simplified model for TDBG with no moiré coupling at the twist interface. By studying the response in this system, we have found that the sign reversal originated from the $180^{\circ}$ rotation of one of the bilayer and that the relative sign between the variants is retained even after the effects of the moiré reconstruction of the bands.
@@ Acknowledgements.
Appendix A Results for $\theta=0.4^{\circ}$ and $2.0^{\circ}$
Figure 8: The band structure of (a) AB-AB and (b) AB-BA stacked TDBG with $\theta=0.4^{\circ}$ and $\Delta=$ 50 meV is shown in the left panels. On the centre and right panels, we show the corresponding $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$ plots, plotting frequency $\omega$ and Fermi energy $E_{\rm{F}}$ on the horizontal and vertical axes respectively. The grey arrows indicate the respective charge neutral gaps.
Figure 9: The band structure of (a) AB-AB and (b) AB-BA stacked TDBG with $\theta=2.0^{\circ}$ and $\Delta=$ 50 meV is shown in the left panels. On the centre and right panels, we show the corresponding $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$ plots, plotting frequency $\omega$ and Fermi energy $E_{\rm{F}}$ on the horizontal and vertical axes respectively. The grey arrows indicate the respective charge neutral gaps.
Here, we show the band structures and shift current plots for AB-AB and AB-BA TDBG for $\theta=0.4^{\circ}$ and $2.0^{\circ}$ with $\Delta=$ 50 meV. The results for $\theta=0.4^{\circ}$ and $2.0^{\circ}$ are shown in Fig. 8 and Fig. 9, respectively. We can see that the signal strength is strong at $\theta=0.4^{\circ}$, while it is suppressed at $\theta=2.0^{\circ}$. This is a similar behaviour to what was seen when the vertical bias voltage $\Delta$ was increased, in which, the widened band gap suppressed the signal strength via the $1/\omega^{2}$ factor. We further note that the sign reversal that was seen below the charge neutrality point in the $(\theta,\Delta)=(0.8^{\circ}$, 50 meV) case is overshadowed at $\theta=2.0^{\circ}$, which is mainly due to the fact that the bands are much more dispersive compared to the $\theta=0.8^{\circ}$ case.
Appendix B Transformation of the conductivity tensor in non-moiré TDBG

### v2

@@ III.1 Evolution with twist angle
We show the intrinsic shift current responses for $\theta=0.4^{\circ},0.8^{\circ},1.4^{\circ},2.0^{\circ}$ and $3.0^{\circ}$ in Fig. 2. In each panel of Fig. 2 (a)/(b), we show the band structures and the density plots of $\sigma^{\mu}_{\mu\mu}(\omega;E_{\rm{F}})$ for AB-AB/AB-BA TDBG for each value of $\theta$. We note that the figures only show the non-vanishing component, $\sigma^{x}_{xx}$ for AB-AB and $\sigma^{y}_{yy}$ for AB-BA; the omitted components are suppressed by the out-of-plane rotational symmetry as argued in the previous section. The shift current signals $\sigma^{\mu}_{\mu\mu}(\omega;E_{\rm{F}})$ are given as colour density plots on the right side of the panels with frequency $\omega$ on the horizontal and Fermi level $E_{\rm{F}}$ on the vertical axis. Therefore, each horizontal strip of the density plot corresponds to a $\sigma^{\mu}_{\mu\mu}$ $\omega$ plot for a certain value of $E_{\rm{F}}$. We note that the scale of the vertical axis is adjusted to match that of the band structure plot and the colour scale of the density plots is shared between all plots.
In the band structures, we observe a gradual formation of moiré subbands as the twist angle decreases, where the hyperbolic bands inherited from AB-stacked bilayer graphene are reconstructed into a series of flat bands. Then, inspecting the $\sigma^{\mu}_{\mu\mu}(\omega;E_{\rm{F}})$ density plots, we find that there is an overall trend of the signal strength to significantly enhance as $\theta$ is decreased. There are a couple of reasons behind this. First, the typical energy gap becomes smaller at lower twist angles, amplifying the response through the $1/\omega^{2}$ prefactor in Eq. (6). Second, the moiré coupling weakens in the large-angle limit. In this regime, the low-energy electronic structure approaches that of decoupled AB-stacked bilayer graphene, which possesses inversion symmetry and therefore suppresses the shift current. As a result, the signal is reduced at larger twist angles, where the properties of the individual bilayers dominate.
We also notice that, at small angles $\theta=0.4^{\circ},0.8^{\circ}$, the low frequency signal is not of a single sign but a series of sign flips is observed as the Fermi level is swept. We can understand this in the following manner. The contribution to the low frequency signal is due to the transitions between neighbouring flat bands. As we change $E_{\rm{F}}$ and the band edge is crossed, the contribution to the signal now comes from a different band transition, leading to a sign flip in the shift current. This also explains why we do not see such abrupt sign changes in the high frequency regime where multiple transitions contribute to the signal. When $\theta$ is increased, we begin to lose this feature as the bands become more dispersive. Since this phenomenon is due to the formation of a series of moiré flat bands, the series of sign flips in the signal is a unique feature of moiré systems.
Figure 2: (a) The band structures of AB-AB stacked TDBG and density plots of $\sigma^{x}_{xx}(\omega;E_{\rm{F}})$ signal for various twist angles. Here, we used $\theta=0.4^{\circ},0.8^{\circ},1.4^{\circ},2.0^{\circ}$ and $3.0^{\circ}$. (b) The corresponding plots for the $\sigma^{y}_{yy}(\omega;E_{\rm{F}})$ signal in the AB-BA variant. We note that the $\sigma^{y}_{yy}$/$\sigma^{x}_{xx}$ component vanishes due to the $C_{2x}$/$C_{2y}$ symmetry in the AB-AB/AB-BA variant. The moiré Brillouin zone in shown in the top right corner.
@@ III.2 Effect of vertical bias voltage
Figure 3: The band structures and $\sigma^{x}_{xx}(\omega;E_{\rm{F}})$ density plots of (a) AB-AB and (b) AB-BA stacked TDBG for various strengths of the vertical bias voltage $\Delta$. Here, we used $\Delta=$ 0, 10, 20 and 50 meV. The grey arrows indicate the respective charge neutral gaps.
Figure 4: The band structures and $\sigma^{y}_{yy}(\omega;E_{\rm{F}})$ density plots of (a) AB-AB and (b) AB-BA stacked TDBG for various strengths of the vertical bias voltage $\Delta$. Here, we used $\Delta=$ 0, 10, 20 and 50 meV. The grey arrows indicate the respective charge neutral gaps.
We now move our attention to the evolution of the response as a result of the application of $\Delta$. We begin with the $\sigma^{x}_{xx}(\omega;E_{\rm{F}})$ response at values of $\Delta=$ 10, 20 and 50 meV and show the band structures of AB-AB and AB-BA stacked TDBG in Fig. 3 (a) and (b) respectively.
In the band structures, as $\Delta$ is increased, we find that the bands become well separated and a wide gap opens at the charge neutrality point (CNP). Focussing on the AB-AB configuration, where $\sigma^{x}_{xx}$ is always non-zero, as $\Delta$ is increased, we see that the overall size of the signal is suppressed. This is due to the fact that the bias voltage increases the gap size, which in turn suppresses the signal via the $1/\omega^{2}$ factor. On the other hand, in the AB-BA variant, as soon as the vertical bias voltage begins to break the $C_{2y}$ symmetry, we observe a finite response. From here, the signal continues to grow in magnitude until $\Delta\sim$ 10 meV, from which the signal follows the same pattern with the $1/\omega^{2}$ suppression for stronger $\Delta$.
@@ III.3 Comparison between AB-AB and AB-BA stacked TDBG
If we begin to compare the density plots between the two stacking configurations, we notice that there is a general positive/negative correlation in the sign of the signal above/below the CNP. The negative correlation is illustrated in Fig. 5 where we show a close-up view of the four density plots below the CNP, indicated by the grey dashed boxes in Fig. 3 and 4. Focussing first on the $\sigma^{x}_{xx}$ response shown in Fig. 5 (a) and (b), we have labelled four regions from A to D where a sign reversal can be seen. For example, in the AB-AB variant, region A has a positive signal, while the corresponding signal in the AB-BA variant is negative. A similar pattern is seen for the other three regions. Likewise, in Fig. 5 (c) and (d), where show the density plots for $\sigma^{y}_{yy}$, we find that regions with the same label have opposite signs. We note that this positive/negative correlation is overshadowed as we turn down $\Delta$ and the bands begin to cluster. We also give the shift current density maps for $\theta=2.0^{\circ}$ in Appendix. B, and show that this correlation is seen at other twist angles.
Figure 5: A close-up view of the $\sigma^{x}_{xx}$/$\sigma^{y}_{yy}$ responses in (a)/(c) AB-AB and (b)/(d) AB-BA TDBG shown in Fig. 3 and 4. We have labelled regions where a sign reversal between the two stacking configurations can be seen.
@@ IV Understanding sign reversal in TDBG shift current
Using this Hamiltonian, we show the band structure (left panel) and the corresponding $\sigma^{y}_{yy}(\omega;E_{\rm{F}})$ response (right panel) of BLG in Fig. 6. The band structure computed from the Hamiltonian in Eq. (11) shows a slice through $k_{x}$ at $k_{y}=0$ around the $K_{+}$ ($K_{-}$) point, plotted in the black solid (red dashed) line. The vertical bias voltage allows for a gap-opening of around 50 meV. On the right panel, we show the shift current response as a colour density plot with frequency $\omega$ on the horizontal and Fermi level $E_{\rm{F}}$ on the vertical axis. The numerical evaluation of Eq. (7) was performed with a broadening of $\eta=$ 1 meV. We can initially see that there is very little or no response at frequencies below 50 meV, which is a reflection of the fact that there are no possible optical transitions in this frequency range due to the gap opened via the bias voltage. As we move up to $\omega=$ 50 meV, we start to see a large response when the Fermi level is set within the gap. This sharp response corresponds to the optical transition from the band edge of the valence to that of the conduction band, where a large joint density of states (JDoS) is realised. We also note that the size of the signal is in the range of 10 mA nm $\rm{V}^{-2}$, which is an order of magnitude smaller compared to the signal seen in TDBG. This result suggests that the effect of the moiré reconstruction of the band structure indeed enhances the shift current response. We also note that the sign of the signal will be reversed through a $180^{\circ}$ rotation of the entire system.
Figure 6: The band structure of AB-stacked BLG from the $K_{+}$ ($K_{-}$) is plotted in the solid black (dashed red) lines in the left panel; a vertical bias voltage of 50 meV is applied across the layers. The shift current $\sigma^{y}_{yy}(\omega;E_{\rm{F}})$ is plotted as a density plot in the right panel.
@@
where we have used the fact that $\sigma^{x^{\prime}}_{x^{\prime}x^{\prime}}=0$. The detail of the derivation is described in Appendix. C. We note that, now, both $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$ are finite since $M_{x}$ symmetry is broken as the two BLG sheets are twisted relative to each other.
@@
If we focus on the signs of the $\sigma^{x}_{xx}$ signal on the centre panels, we can see that there is a positive/negative correlation above/below the charge neutrality point between the two variants. This can be understood from the fact that the relative orientation of the BLG sheets are aligned/antialigned in the AB-AB/AB-BA variant. As the shift current is related to the polarisation of the material, the $180^{\circ}$ rotation of the second bilayer leads to a sign reversal of the signal. The same effect is seen in the $\sigma^{y}_{yy}$ signal on the right panels. This qualitatively explains the sign reversal that was seen in the TDBG calculation, and it reveals that this relationship is retained even after the effect of the moiré coupling.
@@
Figure 8: The band structure of (a) AB-AB and (b) AB-BA stacked uncoupled TDBG is shown in the left panels. The plot is around the $K_{+}$ point, the twist angle is set to $\theta=0.8^{\circ}$ and a vertical bias voltage of $\Delta=$ 50 meV is applied. The contribution from the first (second) BLG is plotted in the solid (dotted) black lines. The corresponding moiré reciprocal lattice vector is also labelled. On the centre and right panels, we show the corresponding $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$ plots, plotting frequency $\omega$ and Fermi energy $E_{\rm{F}}$ on the horizontal and vertical axis respectively.
@@ V Conclusion
We have performed a theoretical investigation on the shift current response in AB-AB and AB-BA stacked twisted double bilayer graphene (TDBG). We investigated the evolution of the TDBG shift current response as the twist angle $\theta$, Fermi level $E_{\rm{F}}$ and vertical bias voltage $\Delta$ were varied. We found that there is a strong enhancement of the signal as $\theta$ is reduced, owing from the small gap size in the moiré subband structure. We also found a series of sign flips in the low frequency regime as $E_{\rm{F}}$ is swept, which is a consequence of the formation of a series of moiré flat bands.
Concerning the effects of the vertical bias voltage, we find that the $\sigma^{\mu}_{\mu\mu}$ component, which vanishes at $\Delta=0$, rapidly increases as a finite $\Delta$ is applied, breaking the out-of-plane rotational symmetry. Importantly, we have found a positive/negative correlation in the sign of the signal between the AB-AB and AB-BA variants above/below the charge neutrality point at large values of the vertical bias voltage. To understand the origin of the sign relationship between the two variants, we considered an uncoupled TDBG model with no moiré interaction at the twist interface. The analysis of this model revealed that the sign reversal arises from the $180^{\circ}$ rotation of one of the bilayers. Remarkably, this relative sign difference between the variants persists even after including the effects of moiré-induced band reconstruction.
@@ Acknowledgements.
Appendix A Deriving the expressions of the shift current response
Here, we give an outline of the derivations of Eq.(6) and Eq.(7). Starting with the former, we refer to the general expression for the second order conductivity tensor given in Ref. [50] Eq.(43), which reads
$\sigma^{\mu}_{\alpha\beta}(\omega_{\Sigma};\omega_{\alpha},\omega_{\beta})$ $=-\frac{q^{3}}{\omega_{\alpha}\omega_{\beta}}\int\frac{d^{2}{\bf\it k}}{(2\pi)^{2}}\sum_{a,b,c}\left[\frac{1}{2}f_{a}\,v^{\mu\alpha\beta}_{aa}+f_{ab}\frac{v^{\mu\beta}_{ab}v^{\alpha}_{ba}}{i\omega_{\alpha}-\varepsilon_{ba}}+\frac{1}{2}f_{ab}\frac{v^{\mu}_{ab}v^{\alpha\beta}_{ba}}{i\omega_{\Sigma}-\varepsilon_{ba}}\right.$
$\ \ \ \ \left.+\frac{v^{\mu}_{ac}v^{\beta}_{cb}v^{\alpha}_{ba}}{i\omega_{\Sigma}-\varepsilon_{ca}}\left(\frac{f_{ab}}{i\omega_{\alpha}-\varepsilon_{ba}}-\frac{f_{bc}}{i\omega_{\beta}-\varepsilon_{cb}}\right)\right]+[(\alpha,i\omega_{\alpha})\leftrightarrow(\beta,i\omega_{\beta})],$ (14)
where $v^{\alpha_{1}\dots\alpha_{n}}=\frac{1}{\hbar^{n}}\frac{\partial^{n}H}{\partial k_{\alpha_{1}}\dots\partial k_{\alpha_{n}}}$ are the generalised $n$-th order velocity operators, $i\omega_{\alpha}$, $i\omega_{\beta}$ are Matsubara frequencies with $i\omega_{\Sigma}=i\omega_{\alpha}+i\omega_{\beta}$ and $[(\alpha,i\omega_{\alpha})\leftrightarrow(\beta,i\omega_{\beta})]$ indicates a symmetrisation between the two indices. The shift current is considered in the case where $\alpha=\beta$ and $i\omega_{\alpha}=-i\omega_{\beta}=i\omega$. Then, we notice that the first and third terms in the square brackets are off-resonant, thus, the contributions to the shift current are given by the second and fourth terms. Focussing on the integrand of the final term with its symmetrised counterpart $\overline{\sigma}$,
$\overline{\sigma}=\sum_{\begin{subarray}{c}a,b,c\\ c\neq a\end{subarray}}\frac{v^{\mu}_{ac}v^{\alpha}_{cb}v^{\alpha}_{ba}}{\varepsilon_{ac}}\left(\frac{f_{ab}}{\hbar\omega-\varepsilon_{ba}+i\eta}-\frac{f_{bc}}{-\hbar\omega-\varepsilon_{cb}+i\eta}+\frac{f_{ab}}{-\hbar\omega-\varepsilon_{ba}+i\eta}-\frac{f_{bc}}{\hbar\omega-\varepsilon_{cb}+i\eta}\right),$ (15)
where we have analytically continued $i\omega\to\hbar\omega+i\eta$, $\eta$ being some infinitesimal quantity. By appropriate exchange of indices, we obtain
$\overline{\sigma}=$ $\hskip 9.6739pt\sum_{\begin{subarray}{c}a,b,c\\ c\neq a\end{subarray}}\frac{v^{\mu}_{ac}v^{\alpha}_{cb}v^{\alpha}_{ba}}{\varepsilon_{ac}}\frac{f_{ab}}{\hbar\omega-\varepsilon_{ba}+i\eta}+\sum_{\begin{subarray}{c}a,b,c\\ c\neq a\end{subarray}}\frac{v^{\mu}_{ca}v^{\alpha}_{ab}v^{\alpha}_{bc}}{\varepsilon_{ac}}\frac{f_{ab}}{\hbar\omega-\varepsilon_{ba}-i\eta}$
$+\sum_{\begin{subarray}{c}a,b,c\\ c\neq b\end{subarray}}\frac{v^{\mu}_{bc}v^{\alpha}_{ca}v^{\alpha}_{ab}}{\varepsilon_{bc}}\frac{f_{ab}}{\hbar\omega-\varepsilon_{ba}-i\eta}+\sum_{\begin{subarray}{c}a,b,c\\ c\neq b\end{subarray}}\frac{v^{\mu}_{cb}v^{\alpha}_{ba}v^{\alpha}_{ac}}{\varepsilon_{bc}}\frac{f_{ab}}{\hbar\omega-\varepsilon_{ba}+i\eta},$ (16)
where $a\leftrightarrow c$, $a\leftrightarrow b$, and $a\to c\to b\to a$ are performed in the second, third, and fourth terms respectively. We can now use
$\frac{1}{x\pm i\eta}=\frac{\mathcal{P}}{x}\mp i\pi\delta(x),$ (17)
on (16) and extract the resonant part to find
$\overline{\sigma}$ $=i\pi\sum_{a,b}f_{ab}\left[\sum_{c\neq a}\frac{-v^{\mu}_{ac}v^{\alpha}_{cb}v^{\alpha}_{ba}+v^{\mu}_{ca}v^{\alpha}_{ab}v^{\alpha}_{bc}}{\varepsilon_{ac}}+\sum_{c\neq b}\frac{v^{\mu}_{bc}v^{\alpha}_{ca}v^{\alpha}_{ab}-v^{\mu}_{cb}v^{\alpha}_{ba}v^{\alpha}_{ac}}{\varepsilon_{bc}}\right]\delta(\hbar\omega-\varepsilon_{ba}).$ (18)
By performing similar operations on the second term in Eq. (14), with the aid of the Hellmann-Feynman theorem
$A^{\alpha}_{ab}=\frac{i\hbar}{\varepsilon_{ba}}v^{\alpha}_{ab},$ (19)
one can reach the form of the shift current given in Eq. (6).
We now move our attention to Eq. (7). By inspecting Eq. (14), we notice that only $\overline{\sigma}$ contributes in the present situation, where our Hamiltonian is linear in ${\bf\it k}$, giving us
$\sigma^{\mu}_{\alpha\alpha}(0;\omega,-\omega)$ $=\frac{i\pi q^{3}}{\omega^{2}}\int\frac{d^{2}{\bf\it k}}{(2\pi)^{2}}\sum_{a,b}f_{ab}\left[\sum_{c\neq a}\frac{-v^{\mu}_{ac}v^{\alpha}_{cb}v^{\alpha}_{ba}+v^{\mu}_{ca}v^{\alpha}_{ab}v^{\alpha}_{bc}}{\varepsilon_{ac}}+\sum_{c\neq b}\frac{v^{\mu}_{bc}v^{\alpha}_{ca}v^{\alpha}_{ab}-v^{\mu}_{cb}v^{\alpha}_{ba}v^{\alpha}_{ac}}{\varepsilon_{bc}}\right]\delta(\hbar\omega-\varepsilon_{ba}).$ (20)
If we focus on the first term in the square brackets, we find that
$-v^{\mu}_{ac}v^{\alpha}_{cb}v^{\alpha}_{ba}+v^{\mu}_{ca}v^{\alpha}_{ab}v^{\alpha}_{bc}$ $=-v^{\mu}_{ac}v^{\alpha}_{cb}v^{\alpha}_{ba}+\left(v^{\mu}_{ac}v^{\alpha}_{cb}v^{\alpha}_{ba}\right)^{*}$
[…]
We can perform the same operation on the second term which brings us to
[…]
where Eq. (7) can be reached by substituting for $q=-e$ and $\omega=\varepsilon_{ba}/\hbar$.
Appendix B Shift current at ($\theta$, $\Delta$) = ($2.0^{\circ},20~{}\rm{meV})$
Figure 9: The band structures and $\sigma^{x}_{xx}(\omega;E_{\rm{F}})$ density plots of (a) AB-AB and (b) AB-BA stacked TDBG. Here, we used ($\theta$, $\Delta)$=($2.0^{\circ}$, 20 meV). In the region indicated by the grey dashed box, we find that the sign of the signal is reversed between the two variants.
In this appendix, we show the AB-AB and AB-BA TDBG band structures and the $\sigma^{x}_{xx}(\omega;E_{\rm{F}})$ density plots at ($\theta$, $\Delta$) = ($2.0^{\circ}$, 20meV) in Fig. 9 (a) and (b), respectively. The band structures for both TDBGs are slightly modified through the application of $\Delta$, however, the signal density plot in AB-AB TDBG remains relatively unchanged compared to the $\Delta=0$ case shown in Fig. 2 (a). For the AB-BA variant, where the signal vanished at $\Delta=0$, we now see a finite signal from the $C_{2y}$ symmetry-breaking provided by the vertical bias voltage.
Focussing on the region indicated by the grey dashed box, we can see that the sign of signal is reversed between the two TDBG stacking configurations. This result provides additional support that the relative sign of the signal is inherited from the uncoupled TDBG, and that it can be seen in a range of twist angles. We further comment that, as we increase the twist angle, the required value of $\Delta$ for the correlation to become visible becomes smaller, thus, easier to observe. This can be understood in the following manner: For the sign reversal to be observed, the low frequency contributions to the shift current must only come from transitions between neighbouring bands, which can be achieved when the bands are well separated. Otherwise, when the bands are clustered, the resulting signal will be a sum of the contributions from several different transitions. Then, cancellations between the contributions will ruin the correlation between the two variants. In the case of $\theta=0.8^{\circ}$, a large $\Delta$ was required to separate the moiré subbands and to observe the negative correlation of the TDBG signals. However, as the twist angle is increased, the bands gain more dispersion and become well separated, owing from the weakening of the moiré coupling. Therefore, unlike the case for $\theta=0.8^{\circ}$, the role of $\Delta$ is solely to provide the symmetry breaking for a finite $\sigma^{x}_{xx}$ in AB-BA TDBG, and thus, a small $\Delta$ is enough to observe the sign reversal.
Appendix C Transformation of the conductivity tensor in uncoupled TDBG